# Preserving the Hermiticity of the One-Body Density Matrix for a Non-Interacting Fermi Gas

### v2

@@ B.3 Gradient identities
$\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}$ $=\frac{1}{s}\sum_{i=1}^{d}\sum_{j=1}^{d}s_{i}\frac{\partial}{\partial r_{i}}\bigg{[}\frac{1}{s}s_{j}\bigg{(}\frac{1}{2}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{j}}+\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{j}}\bigg{)}\Bigg{]}$
$=\frac{1}{s^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}s_{i}\Bigg{[}\frac{1}{2}\bigg{(}\delta_{ij}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{j}}+\frac{s_{j}}{2}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{i}\partial R_{j}}$
$\quad+s_{j}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{i}\partial R_{j}}-\frac{s_{i}s_{j}}{s^{2}}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{j}}\bigg{)}$
$\quad+\bigg{(}\delta_{ij}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{j}}+\frac{s_{j}}{2}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{i}\partial s_{j}}$
$\quad+s_{j}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{i}\partial s_{j}}-\frac{s_{i}s_{j}}{s^{2}}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{j}}\bigg{)}\Bigg{]}$
$=\frac{1}{s^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}\bigg{(}\delta_{ij}\frac{\partial V(\mathbf{R})}{\partial R_{j}}+s_{j}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{i}\partial R_{j}}-\frac{s_{i}s_{j}}{s^{2}}\frac{\partial V(\mathbf{R})}{\partial R_{j}}\bigg{)}$
$=\frac{1}{s^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}s_{i}s_{j}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{i}\partial R_{j}}$
$=\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s},$
where we applied equation 2.5 in line one, and $\delta_{ij}$ is a Kronecker delta: $\delta_{ij}=1$ when $i=j$ and $\delta_{ij}=0$ when $i\neq j$. In between lines two and four we used the fact that both $\partial V(\mathbf{R}+\mathbf{s}/2)/\partial R_{i}=\big{(}\partial(R_{i}+s_{i}/2)/\partial R_{i}\big{)}\partial V(\mathbf{R}+\mathbf{s}/2)/\partial(R_{i}+s_{i}/2)=\partial V(\mathbf{R})/\partial R_{i}$ and $\partial V(\mathbf{R}+\mathbf{s}/2)/\partial s_{i}=\big{(}\partial(R_{i}+s_{i}/2)/\partial s_{i}\big{)}\partial V(\mathbf{R}+\mathbf{s}/2)/\partial(R_{i}+s_{i}/2)=(1/2)\partial V(\mathbf{R})/\partial R_{i}$, where $i=1,2,...,d$ and $s=\sqrt{\mathbf{s}\cdot\mathbf{s}}$ with $\mathbf{s}=\mathbf{r}-\mathbf{r^{\prime}}$.

### v3

@@ C.2 Symmetric WKODM expansion
$-\frac{\beta^{2}}{2m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}+\frac{\beta}{2m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{i}{{m}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})\tilde{w}_{1}+\frac{{\partial}\tilde{w}_{2}}{{\partial\beta}}\bigg{)}=0.$ (3.7)
$\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}{\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\cdot\mathbf{p}}\big{)},$ (3.8)
$\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=-\frac{\beta^{{2}}}{4m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}$
$\quad{-\frac{\beta^{4}}{8m^{2}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\cdot\mathbf{p}\big{)}^{2}+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})^{2}V(\mathbf{R}+\mathbf{s}/2)}.$ (3.9)
Aside from a minus sign difference in equation 3.8, equations 3.8 and 3.9 are the same in form as their non-symmetric counterparts $w_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ and $w_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ respectively, as shown in reference [10]. To keep our notation consistent with the non-symmetric WKODM method, we utilize the fact that equation 3.4 remains invariant when the integrand undergoes the momentum transformation $\mathbf{p}\rightarrow-\mathbf{p}$ (this transformation does not include changing the sign of the differential element $d^{d}p$). This is shown below by equation 3.10, where $f(\mathbf{p})$ is any function which is either a polynomial or can be expressed as a polynomial by a Taylor expansion.
Applying equation 3.10 allows us to re-express equation 3.4 as
where we have defined $w(\mathbf{R},\mathbf{s},\mathbf{p};\beta)=\tilde{w}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)$. Applying the semiclassical expansion from before, $\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ and equations 3.8 and 3.9, we can approximate equation 3.11 by making the substitution $w(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, where
$w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\tilde{w}_{1}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)=-\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}{\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\cdot\mathbf{p}}\big{)},$ (3.12)
$=-\frac{\beta^{{2}}}{4m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}$
$\quad{-\frac{\beta^{4}}{8m^{2}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\cdot\mathbf{p}\big{)}^{2}+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})^{2}V(\mathbf{R}+\mathbf{s}/2).}$ (3.13)
Equations 3.12 and 3.13 are respectively in the same form as the non-symmetric expressions of $w_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ and $w_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$, given in reference [10]. Similar to the procedure we followed both there and within Section 2 for the KODM derivation (based on reference [10]), we are finally in a position to derive the symmetric WKODM $\rho_{WK}(\mathbf{R},\mathbf{s})$ by utilizing inverse Laplace transforms and separating terms based on their order of $\hbar$. We find to second order
$\rho_{WK}(\mathbf{R},\mathbf{s})$ ${=}B_{0}^{{WK}}+B_{1}^{{WK}}+B_{2}^{{WK}};$ (3.14)
$B_{0}^{{WK}}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}^{{WK}}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.15)
$B_{1}^{{WK}}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}^{{WK}}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.16)
$B_{2}^{{WK}}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}^{{WK}}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.17)
where each term of the approximate symmetric Bloch matrix $C^{{WK}}(\mathbf{R},\mathbf{s};\beta)\approx C_{0}^{{WK}}(\mathbf{R},\mathbf{s};\beta)+C_{1}^{{WK}}(\mathbf{R},\mathbf{s};\beta)+C_{2}^{{WK}}(\mathbf{R},\mathbf{s};\beta)$ is given by
$C_{0}^{{WK}}(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (3.18)
$C_{1}^{{WK}}(\mathbf{R},\mathbf{s};\beta)$ $={\frac{g\hbar}{(2\pi\hbar)^{d}}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.19)
$C_{2}^{{WK}}(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta).$ (3.20)
From hereon, the calculations and approximations we perform to fully simplify and combine equations 3.12, 3.13, 3.14, 3.15, 3.16, 3.17, 3.18, 3.19 and 3.20 are identical to those used in Section 2 for the KODM method, hence to avoid redundancy, we will not explicitly show each step. After simplification one exactly finds that in symmetric coordinates $B_{0}^{{WK}}=A_{0}^{{WK}}$ [equation 2.20], $B_{1}^{{WK}}=A_{1}^{{WK}}$ [equation 2.26], and $B_{2}^{{WK}}=A_{2}^{{WK}}$ [equation 2.27]. Therefore, combining these results together via equation 3.14, we find that the symmetric WKODM is exactly identical to the symmetric KODM [equation 2.28], which we know is exactly the $d$-dimensional GVODM originally proposed in reference [23].